OpenStudy (anonymous):

solve the exponential expression: 11(13-6^5-3)=10

OpenStudy (anonymous):

11(13-6\[11(13-6^{5x-3})=10\]

OpenStudy (anonymous):

answer is 0.878212996. take constants on to one side of the equal sign and keep 6^(5x-3) on the other side. you ll get 6^(5x-3) = 133/11 then take logarithms on both side and then solve for x.